OpenStudy (anonymous):

HELP! There are 6 girls and 5 boys in a science class. The teacher chooses two students to lead a class discussion. What is the probability of selecting a girl, then a boy?

OpenStudy (anonymous):

Possible answers 3/110 3/11 36/121 6/11

OpenStudy (anonymous):

Girls, 1:6, Boy, 1:5, Might be Wrong Or Right.

OpenStudy (anonymous):

Oh wait, girl 6:11, and boy, 3:11

OpenStudy (anonymous):

lol.. might be wrong or right...:)

OpenStudy (kinggeorge):

Since it's asking for the probability of a girl, and THEN boy, that actually makes this easier. We have to find the probability that a girl was chosen first, and then the probability that a boy was chosen after that. The probability of a girl being chosen first, is just \[\frac{6}{11}\]since there are 6 girls, out of 11 people. The probability of choosing a boy after this, is \[\frac{5}{10}=\frac{1}{2}\]because there are 5 boys left, but only 10 students to choose from now since we've already removed one girl. That means the probability you're looking for is just \[\frac{6}{11}\cdot\frac{1}{2}=\frac{6}{22}=\frac{3}{11}\]

OpenStudy (anonymous):

Wait, I'm right??!?! YES I AM EXTREMELY SMART! *5 years later, president of the usa*

OpenStudy (anonymous):

yay!!! i'll vote for you...:)

OpenStudy (anonymous):

yay

OpenStudy (kinggeorge):

I wish there was a facepalm emoticon I could easily draw :P

OpenStudy (anonymous):

i know.

OpenStudy (anonymous):

lol! thanks to all!!! you all are genius! :P

OpenStudy (anonymous):

LIKE A BOSS :P